Circular No. 6872

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0421+560 AND CI CAMELOPARDALIS R. M. Hjellming and A. J. Mioduszewski, National Radio Astronomy

Observatory, report: “VLA images obtained on Apr. 5.08 and 6.94 UT at 22.5 GHz show that the CI Cam radio source had become resolved. The extended emission was just becoming apparent on Apr. 5.08, and by Apr. 6.94 it had the appearance of a roughly symmetrical S-shaped twin-jet, strikingly similar to the radio jets of V1343 Aql = SS 433. The outermost pair of emission components was separated by ∼ 0′′.33 on Apr. 6.94; so if one assumes a beginning at the start of the x-ray outburst on Mar. 31.6, the proper motion of each component is 26 mas/day. Assuming a distance of 1 kpc (Chkhikvadze 1970, as quoted by Bergner et al. 1995, A.Ap. Suppl. 112, 221), this corresponds to apparent tangential velocities of 0.15c. The values of proper motion and apparent tangential velocity may be larger if jet ejection occured after the beginning of the x-ray outburst. Better constraints on these values will be obtained after analysis of a series of images. The presence of relativistic jets from CI Cam supports its identification with XTE J0421+560.”

Y. Ueda, M. Ishida, H. Inoue, and T. Dotani, Institute of Space and Astronautical Science; W. H. G. Lewin, Massachusetts Institute of Technology; and J. Greiner, Astrophysical Institute, Potsdam, report: “XTE J0421+560 was observed with ASCA from Apr. 3.31 through 4.14 UT (net exposure 39 000 s). The source was detected at α = 4^{h}19^{m}41^{s}, δ = +56^{o}00′.2 (equinox 2000; error radius 0′.5), which is consistent with the VLA determination (IAUC 6857). The mean flux of the source was 9 × 10^{−10} erg s^{−1} cm^{−2} (2–10 keV), corresponding to ∼ 40 mCrab. The source flux showed a gradual decrease during the observation (e-folding decay timescale of 32 hr), accompanied by a slight softening of the energy spectrum. We detected a prominent iron emission line at 6.7 keV (equivalent width 0.6 keV) indicating a thermal origin of the x-ray emission. The overall energy spectrum could be reproduced by a two-temperature model (5.5 and 0.7 keV) of a thermal emission from the optically thin plasma.”

CORRIGENDA

IAUC 6868, GRB 980329, line 5: for (see IAUC 6864) read [GRB Coordinates Network Circular No. 40 (GCN 40; URL given on IAUC 6864); see

  also http://gcn.gsfc.nasa.gov/gcn/gcn main.html]

IAUC 6869, S/1997 U 2, line 7, and IAUC 6870, S/1997 U 1, line 11: for D. Graham read D. Groom

1998 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green